OpenStudy (anonymous):

Two students in your class, Hunter and Maggie, are disputing a function. Hunter says that for the function, between x = -2 and x = 2, the average rate of change is 0. Maggie says that for the function, between x = -2 and x = 2, the graph goes up through a turning point, and then back down. Explain how Hunter and Maggie can both be correct, using complete sentences.

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

we already did this one, didn't we?

OpenStudy (anonymous):

uh no!

OpenStudy (anonymous):

oh yeah we did sorry wrong one

OpenStudy (solomonzelman):

it's alright

OpenStudy (solomonzelman):

post the actual question (if any) below, in your next reply.

OpenStudy (anonymous):

In your lab, a substance's temperature has been observed to follow the function T(x) = (x - 2)3 + 8. The turning point of the graph is where the substance changes from a solid to a liquid. Explain to your fellow scientists how to find the turning point of this function, using complete sentences.

OpenStudy (solomonzelman):

okay, a turning point is a change either: was decreasing, and now increasing was increasing, and now decreasing

OpenStudy (anonymous):

ok

OpenStudy (solomonzelman):

the easiest thing would be to graph it, are you allowed to do this?

OpenStudy (anonymous):

no its a written response

OpenStudy (anonymous):

or maybe we can and then explain the graph

OpenStudy (anonymous):

but i cant insert a graph

OpenStudy (solomonzelman):

ok, sure. In this case you will have only one turning point, it is the vertex. in \(\large\color{black}{ S(x)=x^3 }\) it would be (0,0) but your function is shifted: \(\large\color{black}{ S(x)=x^3 ~~~~\Rightarrow~~~T(x)=(x\color{red}{-2})^3\color{blue}{+8} }\) (shifted 2 units \(\large\color{black}{ \color{red}{\rm to~the~right} }\) and 8 units \(\large\color{black}{ \color{blue}{\rm up} }\) )

OpenStudy (solomonzelman):

type if you want to say something, regardless of if I am typing or not.

OpenStudy (anonymous):

ok i will

OpenStudy (solomonzelman):

openstudy is loaded with viruses.. I disconnected 50 times just now

OpenStudy (solomonzelman):

anyway, can you tell me where the turning point is?

OpenStudy (anonymous):

what does that mean again?

OpenStudy (anonymous):

is it -1,-1 and 1,1?

OpenStudy (solomonzelman):

in a case y=x^3 it would be the origin, your function is shifted like this: (0+2 , 0+8 )

OpenStudy (solomonzelman):

so what is the turning point?

OpenStudy (solomonzelman):

I shifted the function 2 units right and 8 units up

OpenStudy (anonymous):

oh smh i was thinking that i just dont like to be wrong

OpenStudy (anonymous):

(2,8)

OpenStudy (solomonzelman):

yes.

OpenStudy (anonymous):

because it started from 0,0 right

OpenStudy (anonymous):

n also (-2,-8)

OpenStudy (solomonzelman):

Yes, and there is a proof from calculus: ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~ \(\large\color{black}{T(x)=(x-2)^3+8 }\) take the derivative: \(\large\color{black}{T'(x)=3(x-2)^2 }\) set T'(x)=0 \(\large\color{black}{0=3(x-2)^2 }\) \(\large\color{black}{0=(x-2)^2 }\) \(\large\color{black}{0=x-2 }\) \(\large\color{black}{2=x }\)

OpenStudy (anonymous):

because its plus or minus right or wrong?

OpenStudy (solomonzelman):

no, just and just 92,8)

OpenStudy (solomonzelman):

oops, (2,8)

OpenStudy (anonymous):

ok thanks

OpenStudy (solomonzelman):

So, 2 is a critical number: \(\large\color{black}{T(\color{red}{2})=(\color{red}{2}-2)^2+8 }\) \(\large\color{black}{T(\color{red}{2})=8 }\)

OpenStudy (anonymous):

yes i see it

OpenStudy (solomonzelman):

yes, you welcome:) (and there is no absolute max or min)

OpenStudy (anonymous):

ok